研究集会「Codes, Lattices, Modular Forms and Vertex Operator Algebras」

標記の研究集会を下記の要領で開催しますので、ご案内申し上げます。

主催者山形大学小関道夫九州大学坂内英一山形大学原田昌晃

日程: 2000 年 10 月 2 日 (月) – 4 日 (水) 場所: 山形大学理学部先端科学実験棟 4 階大講義室

プログラム

10 月 2 日 (月)

10:00–10:10 挨拶 10:10–11:00 Rudolf Scharlau (Universität Dortmund, Germany)

Classifying lattices using modular forms 11:10–12:00 Robin Chapman (University of Exeter, UK)

Double circulant constructions of lattices 13:30–14:20 Ming-Chang Kang (National Taiwan University, Taiwan)

Lüroth’s Theorem revisited 14:30–15:10 Steven T. Dougherty (University of Scranton, USA)

Shadows of codes over rings of order 4 15:30–16:20 Michel Deza (Ecole Normale Supérieure, France)

Rank 1 forms, closed zones and laminae 16:30–17:00 青木宏樹 (立命館大)

ヤコビ形式による保型形式の次元の評価 (懇親会)

10 月 3 日 (火)

9:30–10:20 伊吹山知義 (大阪大)

A general theory of differential operators on Siegel modular forms and special polynomials 10:30–11:20 Boris Venkov (Steklov Mathematical Institute in St. Petersburg, Russia)

On Siegel modular forms of weight 12 11:30–12:00 石田信彦 (大阪工業大非常勤講師)、石井伸郎 (大阪府立大)

Generators and defining equation of the modular function field of the group Γ_{1}(N) 13:30–14:20 宮本雅彦 (筑波大)

VOAs and modular forms of rational number weights 14:30–15:00 Haesuk Kim (Pohang University of Science and Technology, Korea)

Codes over ℤ_{2m} and Jacobi forms of genus n 15:20–16:10 桂田英典 (室蘭工大)

Dirichlet series for Siegel modular forms and quadratic forms 16:20–17:00 北岡良之 (名城大)

３次体の単数の分布について

10 月 4 日 (水)

9:30–10:20 Siegfried Böcherer (Universität Mannheim, Germany)

Modular forms of high vanishing order and applications to quadratic forms 10:30–11:20 Patrick Solé (CNRS, ESSI, France)

Rankin-Cohen brackets and transvectants 11:30–12:00 Jon-Lark Kim (University of Illinois at Chicago, USA)

On additive self-dual codes over GF(4) and lattices 13:20–14:10 山田裕理 (一橋大)

Lattice VOAs and Virasoro algebras 14:20–14:50 北詰正顕 (千葉大)

Classification of 12-dimensional even unimodular Gaussian lattices 15:00–15:20 大浦学 (九州大)

A theta relation in genus 4 15:20–15:40 城本啓介 (熊本大)

Maximum distance codes over finite rings 15:40–16:00 原田昌晃 (山形大)

ℤ_{22}-code construction of the Leech lattice

なおこの研究集会の講演者旅費は、坂内悦子、宗政昭弘及び主催者の科学研究費より援助を受けています。

懇親会： 2 日 (月) に懇親会を予定しております。会場準備の為に参加人数を把握したいので、参加希望の方は 9 月 27 日 (水) までにメイルにて原田 (mharada@sci.kj.yamagata-u.ac.jp) までご連絡下さい。

交通案内：バス：山形駅前バスのりばより「県庁前」行きバスにて、「南高前」バス停下車徒歩 5 分。徒歩：山形駅より徒歩 20 分。

最新案内:

http://ksmalpha.kj.yamagata-u.ac.jp/~harada/Yamagata00/ にこの集会に関する最新案内を

準備していますのでご覧ください。

この研究集会に引き続き 10 月 5 日から 8 日には、蔵王エコーホテル (蔵王温泉 (山形市)) にて、「INTERNATIONAL WORKSHOP ON COMBINATORICS(兼、第 12 回日本−フランス組合せ論ワークッショップ)」が開かれます。